OpenStudy (anonymous):

Is (9, -1) a solution of x+5y=4

OpenStudy (igreen):

To find out input (9, -1) into x + 5y = 4. Where x = 9, and y = -1: \(x + 5y = 4\) \(9 + 5(-1) = 4\) \(9 - 5 = 4\) \(4 = 4\) So \('(9, -1)'\) is a solution to \('x + 5y = 4'\).